12.Timekeeper Inc. manufactures clocks on a highly automated assembly line. Its costing system uses two cost categories, directCosts for June 2017: Work in process, beginning inventory: Direct materials Conversion costs Direct materials costs added dur
Answer #1

Answer

Physical unit

Particulars units
Beginning work in process inventory 390
Started 1040
Total 1,430
Completed and transferred out 1,230
Ending work in process inventory 200
Total 1,430

Equivalent unit

Equivalent unit of conversion cost

= 1,230 + (200 × 85%) = 1,400 units

Cost per equivalent unit of conversion cost

= (138,500 + 403,500) / 1,400 = $387.14

Conversion cost per equivalent unit in June is $387.14

Option D $387.14 is correct option.

The above calculations clearly indicate that other options are incorrect.

If YOU are like the answer please upvote.